# Effective Generative AI: The Human-Algorithm Centaur

### v1

@@ Abstract.
Advanced analytics science methods have enabled combining the power of artificial and human intelligence, creating centaurs that allow superior decision-making. Centaurs are hybrid human-algorithm AI models that combine both formal analytics and human intuition in a symbiotic manner within their learning and reasoning process. We argue that the future of AI development and use in many domains needs to focus on centaurs as opposed to traditional AI approaches. This paradigm shift from traditional AI methods to centaur-based AI methods raises some fundamental questions: How are centaurs different from traditional human-in-the-loop methods? What are the most effective methods for creating centaurs? When should centaurs be used, and when should the lead be given to traditional AI models? Doesn’t the incorporation of human intuition—which at times can be misleading—in centaurs’ decision-making process degrade its performance compared to traditional AI methods? This work aims to address these fundamental questions, focusing on recent advancements in generative AI, and especially in Large Language Models (LLMs), as a main case study to illustrate centaurs’ critical essentiality to future AI endeavors.
@@ 1. Introduction
Developing analytics science methods that can enable combining the power of artificial and human intelligence has brought the concept of centaurs from myth to reality. In the Greek mythology, centaurs are half-human and half-horse creatures. In modern analytics science, they refer to systems that allow superior decision-making by combining the power of both humans and trained algorithms. One of the main users in the U.S. has been the Defense Department, which has been working with tech companies to combine the power of algorithms with the capabilities of humans (1). The concept has attracted the attention of the U.S. military, both in research programs at the Defense Advanced Research Projects Agency and the Pentagon’s third-offset strategy for military advantage (2). Robert O. Work, for example, who was the deputy secretary of defense under Presidents Trump and Barack Obama, advocated for the idea of centaur weapons systems, which would require human control, instead of pure Artificial Intelligence (AI) systems, and could combine the power of AI with the capabilities of humans (3).
@@
In this article, we focus on recent advancements in Generative AI, and especially in Large Language Models (LLMs). We first present a framework that allows an understanding of the core characteristics of centaurs. We argue that the incorporation of human intuition and symbiotic learning—a specific learning method based on a human-algorithm symbiosis that we will define and theorize—are key characteristics of centaurs that distinguish them from traditional AI models. Using these core characteristics, we also present a few specific methods of creating centaurs. We then argue that the growth and success of LLMs are to a great extent due to the fact that they are moved from traditional AI algorithms to human-algorithm centaurs. We present various evidence to demonstrate this, particularly by focusing on the advantages of the so-called “fine-tuning” approaches such as the Reinforcement Learning with Human Feedback (RLHF) method used in various LLMs (e.g., OpenAI’s GPT-4, Anthropic’s Claude, Google’s Bard, and Meta’s LLaMA 2-Chat). We also discuss evidence showing that when such fine-tuning approaches are combined with human preference data, they can turn Generative AI tools into cognitive models, capable of representing human behavior. In addition, we elaborate on the main advantages of centaurs: enhanced interpretability and adaptability to behavioral tasks; removing barriers with respect to algorithm aversion, human aversion, and casual aversion; and better performance on challenging prediction tasks, such as those with fuzzy objectives or low-quality datasets. We then briefly conclude by discussing two main points: (1) recent advancements in creating centaurs have moved us closer to reaching the goals that the founding fathers of AI—John McCarthy, Marvin Minsky, Nathaniel Rochester, and Claude Shannon—stated in 1955 as part of their proposed 2-month, 10-man study of AI to be held at Dartmouth; and (2) the future of AI development and use in many domains will most likely need to focus on centaurs as opposed to other traditional AI approaches.
@@ 2. Motivation
Nonetheless, human intuition is often very powerful, especially when we want to make quick decisions. Put it differently, while intuition often misfires when we are dealing with complex problems that require careful analytics (e.g., in predicting whether a particular startup will be successful and is worth large investment (startup), finding ways to reduce incidents of diabetes for organ transplanted patients (10; 11), deciding upon cell formation and layout design for a cellular manufacturing system (12), or finding most effective ways of saving lives in emergency rooms (13; 14)), it can be very useful when using data, models, and careful analytics is not an option. Malcolm Gladwell’s popular book “Blink: The Power of Thinking Without Thinking” provides various examples of this, including when police officers need to quickly decide whether to shoot a suspect (15). Good intuitive decision-making also helps fire fighters when they face a burning building (16).
@@
The complementary functionalities of formal analytics and human intuition are the key to centaurs’ potential of working better than the best algorithms, even in tasks in which human intuition is relatively weak. In our experiments with the Mayo Clinic, for example, the ranking in performance (from best to worst) was as follows: human-algorithm centaur, algorithm, and human experts. Notably, the best human expert’s performance was significantly below that of an AI model that was not benefiting from human intuition. Yet, feeding human intuition to it could create a desirable boost in performance.
@@
2.1. Main Advantages of Centaurs Compared to Traditional AI Models
@@
2. Reduced algorithm aversion: Algorithm aversion has been widely reported as a main hindrance in creating impactful AI in applications in which the final decision-maker is a human (31; 32; 33). Through experiments, we have found that utilizing centaurs can reduce algorithm aversion, mainly because recommendations from centaurs are closer to human way of thinking since centaurs incorporate human intuition (sor). For example, using the “weight on advice measure,” we found that physicians are more willing to consider recommendations that come from centaurs than those from traditional AI models (sor). Furthermore , the recommendations coming out of centaurs represent less “human aversion” in that, not surprisingly, they better represent human behavior and preferences.
@@
6. Enhanced performance on low-quality datasets: A key trait of AI models is the ability to generalize from one dataset, a training set, to other unseen datasets. However, training datasets may contain flaws that inhibit the model’s generalization ability. For instance, the training set at hand can be unrepresentative of the general population. Alternatively, the training dataset may be representative of the general population at train time, $t_{train}$, but not at deployment time, $t_{test}=t_{train}+t^{\prime}$, (i.e. concept drift). Human intuition can be used to mitigate such generalization caveats by adding more nuanced yet general reasoning into the model, enabling it to generalize to unseen, potentially novel, examples (conc1; conc2; conc3; conc4).
@@ 3. Centaur Versus Human-in-the-Loop Approaches
Workload partitioning approaches. In this set of approaches, one typically divides the prediction workload between the human and the machine based on pre-selected criteria. That is, each record is processed either by the machine, the human, or both the machine and the human where the goal is to minimize the number of records that humans are required to process. This approach is oftentimes used in domains in which, on the one hand, experts’ annotation time is expensive, yet, on the other hand, the degree of error tolerance is low. In such case, the minority of “complex” cases will be processed by the human whereas the majority of cases will be processed by the machine. The allocation of records either to the human expert or to the model can be done either based on manually selected criteria such as the risk associated with the record, or based on automatic criteria in which the model itself can decide who should be assigned with the record (e.g., based on its confidence level on the record). Workload partitioning schemes are thus oftentimes deployed in high-risk domains, such as applications related to security and health. While in some works humans and machines share the same workload (work3; work4), in most workload-partitioning works the majority of the workload is put on the AI algorithm, and only the minority of harder or riskier tasks are assigned to humans (work1; work2).
@@
Meta-Learning approaches. In meta-learning approaches, human feedback is neither used for preprocessing actions such as training set organization nor is combined in the learning process. Instead, human feedback is used for different meta-learning tasks such as model evaluation, hyperparameter tuning, and model calibration, among others. Importantly, such feedback is not used to further train the model, but merely to further improve the model without directly combining the feedback in the training process. For example, (me) builds an anomaly-detection ensemble and uses human experts’ opinions on a subset of query records for tuning the ensemble’s hyperparameters: the weight assigned to each model in the ensemble. Likewise, (calibration) presents a human-preference-based calibration approach, suggesting calibration with respect to the human’s confidence in her own predictions produces calibrated estimates that lead to optimal decisions. Finally, (humaneval1; humaneval2) use humans as model evaluators, where humans observe the learned model’s outputs on a representative set of inputs and evaluate its output quality by ranking to what extent it aligns with their subjective preferences.
In all of the approaches discussed above, the human functions mainly as a “service provider:” she is either treated as an oracle, providing labeling services, or as a teacher, providing preprocessing services such as training set design. Notably, the human is not treated as an equal of the learning model who can directly incorporate her own preferences or opinions, such as preference-based labels, as part of the learning process. Specifically, such preference-based human input results from the human’s unique judgment, which is subjective in nature, as opposed to the human’s perception as an oracle where the human input is assumed to always coincide with a certain objective truth (e.g. ground-truth labels). In those human-in-the-loop approaches where humans are not merely used as service providers but are rather queried for their own subjective preferences, such preferences are not combined directly in the learning algorithm. Instead, they are used orthogonally to the AI model, either by working on a different set of records than the machine (workload partitioning approaches) or by providing preference feedback that is not incorporated directly in the learning process (meta-learning approaches).
@@ 4. Symbiotic Learning: A Core Characteristic of Centaurs
As mentioned in previous sections, the distinguishing characteristic of centaurs from traditional human-in-the-loop approaches is symbiotic learning. Symbiotic learning departs from the methods discussed in Section 3 in that it implements a mutual human-AI learning process where the human is not treated merely as an oracle or a preprocessor, but rather as an equal partner in the learning process by incorporating her unique input into the learning process. Such human input includes intuition, subjective preferences, behaviors, or opinions, rather than ground-truth labels. Thus, in summary, the core differences between human-in-the-loop approaches as described in Section 3 and symbiotic learning are two-fold:
1. Symbiotic learning allows for the fact that the human input is often in the form of intuition, a subjective opinion, a behavior, or a preference rather than a ground-truth label.
@@
Symbiotic learning defines a new paradigm, where humans do not function merely as service providers supporting the AI model, but rather as full partners to the model in the learning process. Such partnership can be tuned by the practitioner using flexible constraints, in order to account for a wide array of tasks and applications. A high-level diagram of symbiotic learning is given in Figure 1. In what follows, we present a mathematical framework that further details this representation.
@@
(1) $\hat{\theta}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{x\sim d_{data}}l_{data}(x;\theta)$
(2) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{x\sim d_{human}}l_{human}(x;\theta)$
(3) $\hat{\theta}_{symbiotic}=\gamma_{symbiotic}(\hat{\theta}_{M},~{}\hat{\theta}_{H})=\operatorname*{arg\,min}_{\gamma\in\Gamma,~{}\mathcal{C}(\gamma,\hat{\theta}_{M})\leq c_{1},~{}\mathcal{C}(\gamma,\hat{\theta}_{H})\leq c_{2}}\mathbb{E}_{x\sim d_{human},d_{data}}[l_{comb}(x;\gamma;\hat{\theta}_{M};\hat{\theta}_{H})]$
@@
Given a data distribution $p_{data}$ and a corresponding empirical sample $d_{data}$, we train a base model parametrized by $\hat{\theta}_{M}$ to minimize an arbitrary loss function, $l_{data}$, on $d_{data}$. Given a human-preference distribution $p_{human}$ and a corresponding empirical sample $d_{human}$, we train a human preference model parametrized by $\hat{\theta}_{H}$ to minimize an arbitrary loss function, $l_{human}$, on $d_{human}$. We denote the space of symbiotic-model parameters by $\Gamma$, and model symbiotic learning as minimizing an arbitrary combination loss $l_{comb}$ on either $d_{data}$ or $d_{human}$ w.r.t symbiotic learning parameters $\gamma\in\Gamma$ and subject to optional constraints on either $\hat{\theta}_{M}$, $\hat{\theta}_{H}$, or both. Finally, the learned parameters can be used to learn the functions denoted by $h$ in the last line, where a mapping between the space of inputs to that of outputs occurs. In this setting, loss functions are quite general and include various aspects such as regularizations that is often used in high-dimensional settings to avoid over-fitting.
@@
Distributions $d_{data}$ and $d_{human}$ can also take different forms in different symbiotic learning techniques. Generally speaking, in supervised learning applications of symbiotic learning such as the one discussed above we assume $d_{data}$ to contain pairs of the following form: $(x_{i},y_{i})\in\mathcal{X}_{data}\times\mathcal{Y}_{data}$ and $d_{human}$ to contain pairs of the following form: $(x^{h}_{i},y^{h}_{i})\in\mathcal{X}_{human}\times\mathcal{Y}_{human}$. In some symbiotic learning techniques, both $d_{data}$ and $d_{human}$ must represent the same underlying distribution, while in other techniques they may represent two different distributions. For instance, an underlying assumption of preference-based augmented-covariate-space approaches (see below) is that $d_{data}$ and $d_{human}$ represent the same underlying distribution, where $\mathcal{X}_{data}=\mathcal{X}_{human}$, $\mathcal{Y}_{data}=\mathcal{Y}_{human}$, whereas in human-preference-based supervised fine-tuning approaches (see below) $d_{data}$ and $d_{human}$ can take different distributions. In fact, in the latter case $d_{human}$ is often labeled, where $y^{h}_{i}$ corresponds to human feedback, whereas $d_{data}$ is unlabeled, and thus $\mathcal{Y}_{data}=\emptyset$.
@@
A practical caveat of that approach is that human feedback is needed on each record in $d_{data}$, both in training and in production, and hence the following must hold $|d_{human}|\geq|d_{data}|$, which is impractical in most scenarios. An alternative approach is to collect human preferences on a different, smaller dataset, so that $|d_{human}|<<|d_{data}|$ while $d_{human}$ is highly representative of $d_{data}$. To this end, for for each $i\in d_{data}$ one can search for its k-nearest neighbors in $d_{human}$, and use a combination of their assigned feedback ($y^{h}_{i}$), such as a majority vote, as an additional feature for $i$ in $h_{\gamma}$. However, it is not immediately clear how to choose the best distance metric between samples in $d_{data}$ and samples in $d_{human}$, and the results might be sensitive to the choice made for this metric.

### v4

@@ Abstract.
Advanced analytics science methods have enabled combining the power of artificial and human intelligence, creating centaurs that allow superior decision-making. Centaurs are hybrid human-algorithm models that combine both formal analytics and human intuition in a symbiotic manner within their learning and reasoning process. We argue that the future of AI development and use in many domains needs to focus more on centaurs as opposed to other AI approaches. This paradigm shift towards centaur-based AI methods raises some fundamental questions: How are centaurs different from other human-in-the-loop methods? What are the most effective methods for creating centaurs? When should centaurs be used, and when should the lead be given to pure AI models? Doesn’t the incorporation of human intuition—which at times can be misleading—in centaurs’ decision-making process degrade its performance compared to pure AI methods? This work aims to address these fundamental questions, focusing on recent advancements in generative AI, and especially in Large Language Models (LLMs), as a main case study to illustrate centaurs’ critical essentiality to future AI endeavors.
@@ 1. Introduction
Developing analytics science methods that can enable combining the power of artificial and human intelligence has brought the concept of centaurs from myth to reality. In the Greek mythology, centaurs are half-human and half-horse creatures. In modern analytics science, centaurs represent strong merges between humans and algorithms, referring to hybrid human-algorithm models that combine both formal analytics and human intuition in a symbiotic manner. One of the main users in the U.S. has been the Defense Department, which has been working with tech companies to combine the power of algorithms with the capabilities of humans (1). The concept has attracted the attention of the U.S. military, both in research programs at the Defense Advanced Research Projects Agency and the Pentagon’s third-offset strategy for military advantage (2). Robert O. Work, for example, who was the deputy secretary of defense under Presidents Trump and Barack Obama, advocated for the idea of centaur weapons systems, which would require human control, instead of pure Artificial Intelligence (AI) systems, and could combine the power of AI with the capabilities of humans (3).
@@
In this article, we focus on the role of centaurs in recent advancements in Generative AI, and especially in Large Language Models (LLMs). We first present a framework that allows an understanding of the core characteristics of centaurs. We argue that the incorporation of human intuition and symbiotic learning—a specific learning method based on a human-algorithm symbiosis that we will define and theorize—are key characteristics of centaurs that distinguish them from other AI models. Using these core characteristics, we also present a few specific methods of creating centaurs. We then argue that the growth and success of LLMs are to a great extent due to the fact that they are moved from pure AI algorithms to human-algorithm centaurs. We present various pieces of evidence to demonstrate this, including the advantages of the so-called “fine-tuning” approaches such as the Reinforcement Learning with Human Feedback (RLHF) method used in various LLMs.
We also discuss evidence showing that symbiotic learning on human intuition data can turn Generative AI tools into cognitive models capable of representing human behavior. We also discuss models in which only some of the conditions for creating centaurs are fulfilled, including In-Context Learning or Chain-of-Thought reasoning methods. We conclude by discussing two main points: (1) recent advancements in creating centaurs have moved us closer to reaching the goals that the founding fathers of AI—John McCarthy, Marvin Minsky, Nathaniel Rochester, and Claude Shannon—stated in 1955 as part of their proposed 2-month, 10-man study of AI to be held at Dartmouth; and (2) the future of AI development and use in many domains will most likely need to focus on centaurs as opposed to other AI approaches.
2. Key Ideas
While the concept of centaurs might sound similar to other hybrid paradigms facilitating collaboration between humans and machines such as Human-In-The-Loop-based (HITP) methods, centaurs propose an inherently different Human-Machine Interaction (HMI) paradigm that creates a merge between humans and algorithms as opposed to just a simple partnership. Similar to the fact that two businesses might either create a partnership or merge, humans and algorithms can either work alongside each other through a simple partnership (while being two separate entities that may or may not listen to one another) as suggested by HITP-type methods, or combine their powers through a strong merge, as suggested by centaurs. The unique characteristics of centaurs that allow them to facilitate such strong merge-based collaboration is the use of symbiotic learning and human intuition data. That is, their strength is in both how they learn, and what they learn, and the existence of both components is critical for creating centaurs. Section 5 provides an elaborated definition of centaurs and symbiotic learning, and presents a mathematical framework of centaurs.
Given that the two core components of centaurs are symbiotic learning and learning from human intuition data, a key question that should be posed is whether intuition is at all useful for learning meaningful generalizations. After all, intuition might seem the opposite of rigorous, or logical, or formal (otte). The consensus of most works is that while statistical rules indeed lead to more accurate results on unconstrained decision-making problems, intuition can surprisingly yield better results in some constrained decision-making problems, where typical constraints include incomplete data, uncertainty, and limited decision time, among others (otte; fastslow; strategic; hbrnew). In data-oriented terms, such constraints may refer to insufficient training instances or the lack of representative training cases, missing values for some of the features in the test case, and limited learning time or inference period. Centaurs, in essence, provide a strong merge that benefits from the complementary effect of intuition-based human decision-making and algorithmic decision-making (sor1; sor), providing a single entity that performs well in both constrained and unconstrained situations. Another scenario where human intuition, and specifically, training a model on human intuition data can be useful is when the loss function used prioritizes decisions that coincide with a human-like decision-making mechanism, even if that does not coincide with the optimal decision to take in the environment. We elaborate on specific cases in which the imitation of human decision-making can be beneficial in Section 7.
Understanding that benefiting from intuition can indeed be useful, the next question is, when can it be trusted for making high-quality decisions? Simon and Kahneman define intuition succinctly as unconscious pattern recognition (simone; fastslow). Kahneman describes the recognition-primed decision model as follows: When faced with a problem in a domain of expertise, we initially do not try to come up with several possible courses of action and then evaluate which one is the best. Instead, our associative memory automatically draws on a collection of patterns in our subconscious mind and comes up with the closest one (fastslow). Based on this definition, Hogarth (hogarth), followed by Kahneman and Klein (kk; klein) suggested that intuitive decisions can only be trusted if they reflect a rich and repeated pattern-matching-like experience in environments with reliable feedback. That is, the environment should be stable enough so that past patterns can still be applicable to new ones. Furthermore, meaningful feedback should be constantly provided, and a rich enough set of stored patterns as well as efficient pattern-matching methods should be readily available.
Through symbiotic learning, centaurs ensure the validity of the use of intuition, as the above works suggest. By choosing a model trained on a large set of patterns and with good-enough pattern recognition abilities, the continuous mutual learning process between the human and the machine provides meaningful feedback to both sides. And the fact that centaurs limit the extent to which intuition can influence the model’s weights, mean that they effectively utilize human intuition only to the extent that is beneficial.
The former becomes even stronger when the model used for pattern-matching is a generative model, including the modern architectures of LLMs as we thoroughly elaborate in Section 6. Traditionally, discriminative models were considered preferable to generative models on pattern recognition tasks (generativebook). While generative models were still considered preferable in cases where training sets are small, fast and accurate inference algorithms for complex generative models have been the bottleneck for the application of such models to large-scale instances (generativebook). A major step towards more effective inference in generative modeling has been the Transformer architecture (transformer) and its application to LLLMs. Notably, recent works have demonstrated impressive results obtained by LLMs on a variety of pattern-recognition problems (pattern; pattern2). Apart from the high effectiveness of LLMs for pattern recognition tasks, an additional benefit that makes LLMs highly suitable to be used as centaurs is the use of methods such as Reinforcement Learning with Human Feedback (RLHF) (rlhforiginal). These methods allow reliable human-based feedback in the form of shaped rewards while guaranteeing the environment remains stable enough by constraining the differences between two consecutive policies.
While modern Generative AI constructs such as LLMs incorporate a built-in collaboration between humans and algorithms in the form of prompting and In-Context Learning, we note that it is unclear to what extent such collaborations have a true effect on model learning. In fact, existing literature (prompt1; prompt2) suggests that prompting does not result in real learning, but rather functions merely as a sophisticated “query normalization” method for building better encoding of user queries. That is in contrast to one of symbiotic learning’s inherent traits, the human input having a direct influence on the model’s weights. We stress that this is only one view of prompting, and understanding the exact influence of prompting on LLMs’ learning and inference is still a matter of debate.
Newer models, such as the newest release of OpenAI’s o1 (o12) take one step forward and incorporate human-machine collaboration not only via ad-hoc prompting, but also as an inherent part of the model’s architecture in the form of human-produced Chains-of-Thoughts (CoTs). While CoTs have been commonly used in an In-Context Learning setting for enriching prompts (CoT Prompting), o1 incorporates such human-produced CoTs within its learning and reasoning process using Reinforcement Learning (RL), where the overall objective is for the model to understand not only what a good outcome is, but also what a good “reasoning process" is. The incorporation of human-produced CoTs within the model’s learning and reasoning process can certainly be seen as one step forward toward creating centaur models; however, we argue that such models are still different than centaurs and, hence, do not provide the strong human-algorithm merge we introduced earlier. The key difference lies in the definition of intuition as suggested by (simone; fastslow), and specifically on its inherent trait of being “unconscious.” According to (klein), this fact can be attributed to the reliance of intuition on tacit knowledge: knowledge gained through experience. Tacit knowledge lacks structured order, is difficult to explain, and often unavailable to consciousness (klein; fastslow; simone). CoT prompting or learning represents the opposite: it is a structured, conscious, and purposely explainable human input that aims at conveying purely logical arguments. Models trained on CoT human input alone, thus, might not be considered full centaurs, since they lack human intuition.
In closing this section, we note that our aim in this paper is mainly (a) starting to raise important questions about human-algorithm centaurs, and (b) providing the first steps for encouraging a wide-ranging debate about their potential in making future AI models more effective. The definitions and arguments we provide, thus, are not black-and-white. Furthermore, they might change over time, especially since it is up to future scholars to create highly effective centaur-based AI models.
The rest of this document is organized as follows. Section 3 presents the motivation for creating and using centaurs. Section 4 characterizes common existing methods for human-machine collaboration, clarifying why centaurs are different. Section 5 presents the notion of symbiotic learning and provides a concrete mathematical framework for creating centaurs. Section 6 discusses why LLMs are highly suitable as base models for creating centaurs, and presents evidence suggesting that centaur-based LLMs can function, at least partially, as cognitive models. Section 7 discusses the advantages and disadvantages of centaur models, their potential use cases, and some ethical considerations. Finally, Section 8 concludes by discussing why future use of AI in various applications will likely be centaur-based.
@@ 3. Motivation
Nonetheless, human intuition is often very powerful, especially when we want to make quick decisions. Put it differently, while intuition often misfires when we are dealing with complex problems that require careful analytics (e.g., in predicting whether a particular startup will be successful and is worth large investment (startup), finding ways to reduce incidents of diabetes for organ transplanted patients (10; 11), deciding upon cell formation and layout design for a cellular manufacturing system (12), or finding most effective ways of saving lives in emergency rooms (sor2; 13; 14)), it can be very useful when using data, models, and careful analytics is not an option. Malcolm Gladwell’s popular book “Blink: The Power of Thinking Without Thinking” provides various examples of this, including when police officers need to quickly decide whether to shoot a suspect (15). Good intuitive decision-making also helps fire fighters when they face a burning building (16).
@@
The complementary functionalities of formal analytics and human intuition are the key to centaurs’ potential of working better than the best algorithms, even in tasks in which human intuition is relatively weak. In the first author’s experiments with the Mayo Clinic, for example, the ranking in performance (from best to worst) was as follows: human-algorithm centaur, algorithm, and human experts. Notably, the best human expert’s performance was significantly below that of an AI model that was not benefiting from human intuition. Yet, feeding human intuition to it could create a desirable boost in performance (sor; sor1).
@@
3.1. Main Advantages of Centaurs Compared to Other AI Models
@@
2. Reduced algorithm aversion: Algorithm aversion has been widely reported as a main hindrance in creating impactful AI in applications in which the final decision-maker is a human (31; 32; 33). Through experiments, we have found that utilizing centaurs can reduce algorithm aversion, mainly because recommendations from centaurs are closer to human way of thinking since centaurs incorporate human intuition (sor). For example, using the “weight on advice measure,” we found that physicians are more willing to consider recommendations that come from centaurs than those from pure AI models (sor). Furthermore , the recommendations coming out of centaurs represent less “human aversion” in that, not surprisingly, they better represent human behavior and preferences.
@@
6. Enhanced performance on low-quality datasets: A key trait of AI models is the ability to generalize from one dataset, a training set, to other unseen datasets. However, training datasets may contain flaws that inhibit the model’s generalization ability. For instance, the training set at hand can be unrepresentative of the general population. Alternatively, the training dataset may be representative of the general population at train time, $t_{train}$, but not at deployment time, $t_{test}=t_{train}+t^{\prime}$, (i.e. concept drift). Human intuition can be used to mitigate such generalization caveats by adding more nuanced yet general reasoning into the model, enabling it to generalize to unseen, potentially novel, examples (conc1; conc2; conc3; conc4; conc5).
@@ 4. Centaur Versus Human-in-the-Loop Approaches
Workload partitioning approaches. In this set of approaches, one typically divides the prediction workload between the human and the machine based on pre-selected criteria. That is, each record is processed either by the machine, the human, or both the machine and the human where the goal is to minimize the number of records that humans are required to process. This approach is oftentimes used in domains in which, on the one hand, experts’ annotation time is expensive, yet, on the other hand, the degree of error tolerance is low. In such case, the minority of “complex” cases will be processed by the human whereas the majority of cases will be processed by the machine. The allocation of records either to the human expert or to the model can be done either based on manually selected criteria such as the risk associated with the record, or based on automatic criteria in which the model itself can decide who should be assigned with the record (e.g., based on its confidence level on the record). Workload partitioning schemes are thus oftentimes deployed in high-risk domains, such as applications related to security and health. While in some works humans and machines share the same workload (work3; work4; refprofasked), in most workload-partitioning works the majority of the workload is put on the AI algorithm, and only the minority of harder or riskier tasks are assigned to humans (work1; work2).
@@
Meta-Learning approaches. In meta-learning approaches, human feedback is neither used for preprocessing actions such as training set organization nor is combined in the learning process. Instead, human feedback is used for different meta-learning tasks such as model evaluation, hyperparameter tuning, and model calibration, among others. Importantly, such feedback is not used to further train the model, but merely to further improve the model without directly combining the feedback in the training process. For example, (anomalyensemble) builds an anomaly-detection ensemble and uses human experts’ opinions on a subset of query records for tuning the ensemble’s hyperparameters: the weight assigned to each model in the ensemble. Likewise, (calibration) presents a human-preference-based calibration approach, suggesting calibration with respect to the human’s confidence in her own predictions produces calibrated estimates that lead to optimal decisions. Finally, (humaneval1; humaneval2) use humans as model evaluators, where humans observe the learned model’s outputs on a representative set of inputs and evaluate its output quality by ranking to what extent it aligns with their subjective preferences.
Other approaches. More recent methods for human-machine collaboration such as Human-Guided Rewards, Human-based Prompt Engineering, Chain-of-Thought Reasoning, Supervised Fine-Tuning on Human Intuition Data, and more, are discussed in Section 5 and Section 6.
In almost all of the approaches discussed above, the human functions mainly as a “service provider:” she is either treated as an oracle, providing labeling services, or as a teacher, providing preprocessing services such as training set design. Notably, the human is not treated as an equal of the learning model who can directly incorporate her own intuition—in the observed form of preferences, actions, decisions, or opinions—as part of the learning process. Such observed signals of human intuition result from the human’s unique judgment, which is subjective in nature, as opposed to the human’s perception as an oracle where the human input is assumed to always coincide with a certain objective truth (e.g. ground-truth labels). In those human-in-the-loop approaches where humans are not merely used as service providers but are rather queried for their own subjective input, such inputs are not combined directly in the learning algorithm. Instead, they are used orthogonally to the AI model, either by working on a different set of records than the machine (workload partitioning approaches) or by providing subjective feedback that is not incorporated directly in the learning process (meta-learning approaches).
@@ 5. Symbiotic Learning: A Core Characteristic of Centaurs
As mentioned in previous sections, the distinguishing characteristic of centaurs from human-in-the-loop approaches is symbiotic learning. Symbiotic learning departs from the methods discussed in Section 4 in that it implements a mutual human-AI learning process where the human is not treated merely as an oracle or a preprocessor, but rather as an equal partner in the learning process by incorporating her unique input into the learning process. Human input includes observed signals of human intuition such as preferences, behaviors, or opinions, and unlike ground-truth labels, such inputs are inherently subjective. The core differences between human-in-the-loop approaches as described in Section 4 and symbiotic learning are three-fold:
1. Symbiotic learning facilitates multi-input learning, in which learning is performed using multiple datasets corresponding to either the same set of observations or different sets of observations. At least one dataset represents human input, and includes observed signals of human intuition such as subjective preferences, behaviors, opinions, or decisions.
@@
3. The influence of human input data on the final model’s parameters is constrained to ensure appropriate reliance on human intuition given the specific traits of the environment, the degree of human expertise, and feedback quality. The constraints are optimized either as a separate learning task or jointly with $\theta_{symbiotic}$.
Symbiotic learning defines a new paradigm, where humans do not function merely as service providers supporting the AI model, but rather as full partners to the model in the learning process. To account for a wide array of tasks and applications, the nature of such partnership is learned on a case-by-case basis using flexible constraints. A high-level diagram of symbiotic learning is given in Figure 1. In what follows, we present a mathematical framework that further details this representation.
@@
(1) $\hat{\theta}_{M}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{z\sim d_{data}}l_{data}(z;\theta)$
(2) $\hat{\theta}_{H}=\operatorname*{arg\,min}_{\theta\in\Theta}\mathbb{E}_{z\sim d_{human}}l_{human}(z;\theta)$
(3) $\hat{\theta}_{symbiotic}=\gamma_{symbiotic}(\hat{\theta}_{M},~{}\hat{\theta}_{H})=\operatorname*{arg\,min}_{\gamma\in\Gamma,~{}\mathcal{C}(\gamma,\hat{\theta}_{M})\leq c_{1},~{}\mathcal{C}(\gamma,\hat{\theta}_{H})\leq c_{2}}\mathbb{E}_{z\sim d_{human},d_{data}}[l_{comb}(z;\gamma;\hat{\theta}_{M};\hat{\theta}_{H})]$
@@
Given a data distribution $p_{data}$ and a corresponding empirical sample $d_{data}$, we train a base model parametrized by $\hat{\theta}_{M}$ to minimize an arbitrary loss function, $l_{data}$, on $d_{data}$. Given a human-preference distribution $p_{human}$ and a corresponding empirical sample $d_{human}$, we train a human preference model parametrized by $\hat{\theta}_{H}$ to minimize an arbitrary loss function, $l_{human}$, on $d_{human}$. We denote the space of symbiotic-model parameters by $\Gamma$, and model symbiotic learning as minimizing an arbitrary combination loss $l_{comb}$ on either $d_{data}$ or $d_{human}$ w.r.t symbiotic learning parameters $\gamma\in\Gamma$ and subject to constraints, $\mathcal{C}$, on either $\hat{\theta}_{M}$, $\hat{\theta}_{H}$, or both. Finally, the learned parameters can be used to learn the functions denoted by $h$ in the last line, where a mapping between the space of inputs to that of outputs occurs. In this setting, loss functions are quite general and include various aspects such as regularizations that is often used in high-dimensional settings to avoid over-fitting.
@@
Distributions $d_{data}$ and $d_{human}$ can also take different forms in different symbiotic learning techniques. Generally speaking, in supervised learning applications of symbiotic learning such as the one discussed above we assume $d_{data}$ to contain pairs of the following form: $z_{i}=(x_{i},y_{i})\in\mathcal{X}_{data}\times\mathcal{Y}_{data}$ and $d_{human}$ to contain pairs of the following form: $z_{i}=(x^{h}_{i},y^{h}_{i})\in\mathcal{X}_{human}\times\mathcal{Y}_{human}$. In some symbiotic learning techniques, both $d_{data}$ and $d_{human}$ must represent the same underlying distribution, while in other techniques they may represent two different distributions. For instance, an underlying assumption of preference-based augmented-covariate-space approaches (see below) is that $d_{data}$ and $d_{human}$ represent the same underlying distribution, where $\mathcal{X}_{data}=\mathcal{X}_{human}$, $\mathcal{Y}_{data}=\mathcal{Y}_{human}$, whereas in human-preference-based supervised fine-tuning approaches (see below) $d_{data}$ and $d_{human}$ can take different distributions. In fact, in the latter case $d_{human}$ is often labeled, where $y^{h}_{i}$ corresponds to human feedback, whereas $d_{data}$ is unlabeled, and thus $\mathcal{Y}_{data}=\emptyset$.
@@
A practical caveat of that approach is that human feedback is needed on each record in $d_{data}$, both in training and in production, and hence the following must hold $|d_{human}|\geq|d_{data}|$, which is impractical in most scenarios. An alternative approach is to collect human preferences on a different, smaller dataset, so that $|d_{human}|\ll|d_{data}|$ while $d_{human}$ is highly representative of $d_{data}$. To this end, for for each $i\in d_{data}$ one can search for its k-nearest neighbors in $d_{human}$, and use a combination of their assigned feedback ($y^{h}_{i}$), such as a majority vote, as an additional feature for $i$ in $h_{\gamma}$. However, it is not immediately clear how to choose the best distance metric between samples in $d_{data}$ and samples in $d_{human}$, and the results might be sensitive to the choice made for this metric.